# Detecting Spatial Dependence in Transcriptomics Data using Vectorised Persistence Diagrams

### v1

@@ 4.5 Synthetic Data Generation
Figure 2: Simulated spatial patterns with varying effect sizes for each spatial domain as indicated in the legend (row 1). Examples of genes simulated with these spatial signals (row 2).
We create in-silico examples of spatial transcriptomics data using SRTsim [42], an R package for the simulation of spatial patterns. In particular, we randomly sample spatial locations on a square domain and then simulate count data using (zero-inflated) Poisson or negative binomial distributions. The choice between these range of distributions of count data is motivated by the fact that the observed distribution and nature of transcriptomics data is highly debated and varies between sequencing technologies and datasets. Zero-inflation aims to account for the excess zeros often encountered in sparsely sampled gene expression data due to a high dropout probability. Across our simulations we thus consider zero-inflation to be a potential source of noise and assess the robustness of the proposed spatially variable gene detection methods against these technical errors. SRTsim allows us choose the following settings for simulations. Globally, we set the mean parameter ($\mu=1$), the dispersion parameter ($s$) and the zero-proportion parameter ($z$). We then define $N$ spatial domains corresponding to distinct spatial shapes that are assigned an effect size $x_{i}$ resulting in an $x_{i}$-fold increase in mean for the specified spatial domain $i\in{0,1,2,..,N}$. The effect size of the background domain is set to $x_{0}=1$. Count data is then sampled separately for each domain with the thus specified parameters and assigned randomly to locations within the shape. We then simulate four distinct spatial patterns, a gradient, two lines, a set of clusters and a cellring illustrated in Fig. 2 with the annotated effect sizes. Otherwise, if no spatial signal is simulated, gene expression is sampled randomly across the locations.
@@ 5 Results
Across our experiments we consider two main goals: (i) using significance testing to determine spatially dependent features, (ii) rank features by their spatial dependence and identify a fixed number of the top most spatially variable genes. Throughout our empirical experiments we then demonstrate the following results:
  • PH-based approaches perform well at detecting SVGs in terms of the area under the precision recall curve and are more robust to zero-inflation i.e. high proportions of excess zeros in gene expression counts than Moran’s I.
@@
  • PH-based spatial dependence test are less correlated to the total sum of feature values than Moran’s I.
We further explore the use of our PH-based approach to detect SVGs for three real spatial omics datasets, for which we plot and examine the top spatially variable features. We thus demonstrate that persistence-based methods, and Betti curves in particular, offer valid alternatives to existing SVG methods, such as Moran’s I.
@@ 5.1.1 PH-based approaches are robust to zero-inflation
We simulate spatial data as described in Section 4.5 and first vary the zero-inflation parameter $z$ from $0.1$ to $0.99$ keeping the dispersion fixed at $0.3$. For each shape and parameter choice, we then simulate $50$ genes that possess the spatial signal and $50$ genes that are randomly sampled without any spatial effect. Then, across each degree of zero-inflation we summarise the performance of each statistical test and scoring method in the following manner: Fig. 3 reports the area under the precision-recall curve (AUPRC) as a function of the degree of zero-inflation for all methods. This comparison is chosen because it allows us to compare all methods, also the ranking-based sepal method, as a single summary statistic per degree of zero-inflation. An AUPRC of $1$ indicates perfect performance, random guessing is expected to result in an AUPRC of $0.5$, and 00 is the worst possible value.
@@ 5.1.2 Persistence-based testing performs well at detecting SVGs
We observe that all persistence-based methods achieve high AUPRC values very close to $1$ until around $70\%$ of zero-inflation across all pattern. This is followed by a decrease in AUPRC, which remains above $0.6$, even for $99\%$ of excess zeros. Overall, these trends show robustness and expressivity of persistent homology as a summary of spatial patterns even in the presence of high zero-inflation.
@@ 5.1.3 Persistence-based methods perform similarly to one another
We further observe that both Betti curves and total persistence follow very similar trends in Fig. 3 with Betti curves performing better on the gradient pattern and total persistence showing higher scores on the streaks pattern. This finding highlights that both approaches are alike in terms of the information they are summarising. Persistent landscapes meanwhile achieve consistently higher AUPRC values of above $0.7$ for higher degrees of zero-inflation across datasets and thus capture trends in spatial variability in a more robust manner for high degrees of dropout. For this experiment, we thus observe that persistence landscape summarise different aspects of persistent homology than alternative summaries.
5.1.4 Persistence-based methods outperform Moran’s I
Our results thus indicate that persistent homology is more expressive at quantifying spatial signals in simulated expression values under the presence of zero-inflation than Moran’s I. That is we observe that by tracking the evolution of connected components via superlevelset filtrations on spatial graphs, we get summary statistics with higher discriminative performance than Moran’s I, which summarises spatial autocorrelation based on summarising local neighbourhoods and is the most commonly used baseline measure for spatial variability. Indeed, via this comparison in terms of AUPRC, we assess that persistence-based summaries offer a valid alternative to established tools from spatial statistics.
@@ 5.1.5 Persistence detects spatial variability across all simulations
We note that our proposed persistence-based tests do well at detecting spatial variability, in particular for low levels of zero-inflation, as would be expected. However, in comparison, we see that some alternative methods do not follow the same clear trend across datasets. SpatialDE starts decreasing in terms of AUPRC at comparably lower proportions of zero inflation. This difference is especially pronounced for the streaks pattern, where SpatialDE performs notably worse than all alternative methods after $60\%$ zero-inflation. Potentially this is due to this simulation only showing two relatively thin lines that are harder for SpatialDE to detect as spatially variably under the presence of excess zeros. Meanwhile, sepal performs very well for higher values of $z$ between $0.6$ and $0.8$ across datasets, showing higher AUPCRs than persistence landscapes. However, sepal fails at detecting the gradient pattern at low degrees of zero-inflation. For example, sepal receiving an AUPRC of only $0.6$ at $z=0.1$ highlighting that diffusion times are not suitable for detecting this smooth trend in spatial variability. We thus demonstrate disadvantages of both sepal and SpatialDE compared to our persistence-based approach. We conclude that PH-based methods achieve overwhelmingly higher or similar AUPRCs to the baseline method Moran’s I highlighting their utility as robust SVG detection methods.
@@ 5.1.6 PH-based methods are more specific than Moran’s I
Extending our observations beyond AUPRC, we next decide to view spatially variable gene detection in the context of hypothesis testing. Fig. 4 then compares the performance of each statistical test in terms of specificity and sensitivity at a fixed $0.05$ significance threshold for corrected p-values. In agreement with results from Byers et al. [11] we observed that at this fixed threshold, total persistence is either similarly or slightly less powerful i.e. less sensitive than Moran’s I. This is followed by Betti curves, which detects a lower proportion of true positives, in particular on the streaks pattern, as well as persistence landscapes which give even more conservative estimate. However, we also observe that all persistence-based approaches show very high specificity across examples, whereas Moran’s I demonstrates the highest false positive rate of around $5\%$ to $25\%$, even after p-value correction, which is undesirable in practice, when the set of spatially variable genes should be correctly identified.
Our findings thus agree with Byers et al. [11], who find that persistence is less sensitive than Moran’s I. However, Byers et al. [11] further claim that persistence cannot replace Moran’s I, a findings we do not fully support. Arguably, specificity as well as sensitivity is important for interpreting spatial dependency tests in practice. Indeed, we observe that at a default $0.05$ significance cutoff, corrected p-values from PH-based approaches achieve higher specificity than Moran’s I, showing that PH-based results can more useful for tasks that require a low false positive rate and benefit from a more conservative assessment. Given the high AUPRC achieved by PH-based methods, we further assess that choosing a higher significance threshold could be a suitable adjustment to achieve higher sensitivity and identify a larger set of spatially variable features while retaining superior specificity.
@@ 5.1.7 Agreement between methods
Figure 5: Spearman correlation between the rankings given by different spatial variability detection methods across the four spatial patterns from Fig. 2. Generally there is low to medium correlation between methods with the highest agreement being between total persistence and Betti curves. Persistence thus captures distinct characteristics in spatial variability not identified by other approaches.
Further, we compute the Spearman correlation between each method’s ranking and report the results in Fig. 5. For persistence-based approaches we rank genes by their p-values, else we use the values of the scores as defined in Section 4.3 for sepal and Moran’s I. Overall, we observe that total persistence and Betti curves perform most similarly across experiments with correlations between 0.64 and 0.82, which is likely driven by their common roots in accumulating persistence features [34]. Further, we find that there is varying but overall moderate levels of agreement between persistence and Moran’s I ($\rho\leq 0.53$), likely explained by the fact that both are summaries of the same spatial graph and understand spatial variability via modelling clustering behaviour, as summarised by connected components in PH, or autocorrelation, as summarised by differences in neighbourhood scores. Interestingly, while sepal follows a fundamentally different modelling approach, correlation between sepal and the other approaches exist at comparable levels to other entries in the correlation matrix and the sepal score is actually most correlated with Moran’s I on the clusters dataset ($\rho=0.58$). Overall, we thus observe that persistence-based permutation testing captures complementary information not encoded by either Moran’s I, sepal or SpatialDE.
@@ 5.2 PH-based approaches identify the top SVGs
Next, we keep the zero proportion at $z=0.1$ and the dispersion parameter at $s=0.3$, following estimations by Zhu et al. [42] from real data, but vary the effect size, by taking $\hat{x_{i}}\leavevmode\nobreak\ =\leavevmode\nobreak\ max(x_{i}/e,1)$ for some $e$ in $\{6,5,4,3,2,1\}$ for each domain $i$ effectively varying the strength of the spatial pattern compared to the background domain for which $x_{i}=1$. For each parameter choice and spatial shape we then simulate 50 features with spatial signals and 50 completely random patterns. We then investigate the identification of top spatially variable genes for different strengths of spatial effects.
@@ 5.2.1 Correlation with the sum of gene expression counts
As reported in Fig. 6, we explore the correlation between each spatial variability score and the total count per gene / i.e. the sum of feature values. We want an ideal SVG detection method to not just correspond to these total counts, but score genes based solely on the existence of spatial dependencies in gene expression. Across datasets, we see that Moran’s I and sepal show the highest correspondence to the sum of feature values. In comparison, persistence-based ranking achieve consistently lower Spearman correlation of $0.6$ to the number of total counts. Persistence landscapes reaching slightly higher correlation than Betti curves and total persistence. Overall, this rank correlation is similar to the one achieved by SpatialDE. Therefore, the comparison highlights that persistence offers a satisfactory alternative to established SVG tests.
@@ 5.2.2 PH-based approaches correctly identify the top SVGs
Across methods, we then compare the ground truth given by the existence of spatial signal with each spatial variability score. Fig. 6 shows the proportion of truly positive features detected by each method. Ideally, a suitable spatial variability detection method would show a high proportion of true positives amongst the top 300 genes (given 300 out of 600 features show spatial patterns). Notably, Betti curves perform consistently well across examples correctly identifying more than $90\%$ of top SVGs. This is followed by total persistence and persistence landscapes, which performs better on the gradient but worse on the cluster pattern. Interestingly, from these examples we see that while the proportion of genes correctly identified is high across scenarios, there is no strong distinction between either persistence- or non persistence based methods based on these examples. However, sepal and Moran’s I seem to perform slightly worse at identifying the most spatially variable features than persistence-based approaches, which together with their correlation to total counts, indicates that persistence-based testing is preferable in this scenario.
@@
Figure 9: Correlation between the ranking of all genes ranked by each method for the mouse olfactory bulb (top), the breast cancer sample (middle), and the heart data (bottom). Each plot shows the spatial distribution of one gene’s $log(X+2)$ transformed expression values. All genes visualised here are examples of the most spatially variable genes as identified by Betti curves that are also highly variable across cells.

### v2

@@ 4.5 Synthetic Data Generation
Figure 2: Simulated spatial patterns with varying effect sizes for each spatial domain as indicated in the legend (row 1). Examples of simulated genes with these spatial signals (row 2). The simulated patterns are showing a gradient, a cellring, a set of clusters, and two lines from left to right.
We create in-silico examples of spatial transcriptomics data using SRTsim [47], an R package for the simulation of spatial patterns. In particular, we randomly sample spatial locations on a square domain and then simulate count data using (zero-inflated) Poisson or negative binomial distributions. The choice between these range of distributions of count data is motivated by the fact that the observed distribution and nature of transcriptomics data is highly debated and varies between sequencing technologies and datasets. Zero-inflation aims to account for the excess zeroes often encountered in sparsely sampled gene expression data due to a high dropout probability. Across our simulations we thus consider zero-inflation to be a potential source of noise and assess the robustness of the proposed spatially variable gene detection methods against these technical errors. SRTsim allows us to choose the following settings for simulations. Globally, we set the mean parameter ($\mu=1$), the dispersion parameter ($r$), and the zero-proportion parameter ($z$). We then define $N$ spatial domains corresponding to distinct spatial shapes that are assigned an effect size $e_{i}$ resulting in an $e_{i}$-fold increase in mean for the specified spatial domain $i\in{0,1,2,..,N}$. The effect size of the background domain is set to $e_{0}=1$. Count data is then sampled separately for each domain with the thus specified parameters and assigned randomly to locations within the shape. We then simulate four distinct spatial patterns: a gradient, two lines, a set of clusters, and a cellring illustrated in Fig. 2 with the annotated effect sizes. These spatial signals are to mimic distinct spatial tissue structures which could occur in real datasets [47, 16, 6]. We also simulate genes without spatial signal, whose gene expression values are sampled randomly across the spatial locations.
@@ 5 Results
Across our experiments we consider two main goals: (i) using significance testing to determine spatially dependent features, and (ii) ranking features by their spatial dependence and identifying a fixed number of the top most spatially variable genes. Throughout our empirical experiments we make the following observations:
  • PH-based approaches perform well at detecting SVGs in terms of the area under the precision recall curve. They are more robust to zero-inflation, i.e. high proportions of excess zeroes in gene expression counts, than Moran’s I.
@@
  • PH-based spatial dependence tests are less correlated to the total sum of feature values than Moran’s I.
We further explore the use of our PH-based approach to detect SVGs for three real spatial omics datasets, for which we plot and examine the top spatially variable features. We thus demonstrate that persistence-based methods, and Betti curves in particular, offer valid alternatives to existing SVG methods, such as Moran’s I. Table 1 summarises the performance of each SVG detection methods across all our experiments, which we will now elaborate on in more detail.
Table 1: Summary of the ranking of each SVG detection method by their performance across experiments. The lower the rank the better. Persistence-based permutation tests achieve high median ranks.
Ranking of SVG Methods
Experiment Dataset Betti Curves Pers. Landscapes Total Pers. SpatialDE Moran’s I sepal
AUPRC (Fig. 3, Sec. 5.1.1 to 5.1.5) Gradient 2 1 2 5 4 6
Cellring 3 2 3 5 6 1
Clusters 2 1 2 6 4 5
Streaks 3 2 3 6 3 1
Sensitivity (Fig. 4, Sec. 5.1.6) Gradient 3 4 1 5 1 /
Cellring 3 4 1 5 1 /
Clusters 3 4 1 5 1 /
Streaks 3 3 1 5 1 /
Specificity (Fig. 4, Sec. 5.1.6) Gradient 2 1 2 4 5 /
Cellring 1 1 1 4 5 /
Clusters 2 1 2 4 5 /
Streaks 1 1 1 4 5 /
Correlation with total counts (Fig. 6, Sec. 5.2.1) Gradient 2 1 2 4 5 /
Cellring 1 1 1 4 5 /
Clusters 2 1 2 4 5 /
Streaks 1 1 1 4 5 /
Top genes identified (Fig. 6, Sec. 5.2.2) Gradient 2 2 2 2 6 1
Cellring 1 2 1 1 6 5
Clusters 1 2 1 1 6 5
Streaks 1 2 1 1 6 5
Number of SVGs (Fig. 7, Sec. 5.3.2) MOB 2 4 3 4 1 /
Breast 2 5 5 3 1 /
Heart 1 4 2 5 3 /
Correlation with total counts (Fig. 7, Sec. 5.3.3) MOB 5 3 4 1 6 2
Breast 3 2 1 4 6 5
Heart 5 1 2 4 3 6
Median Rank 2 2 2 4 4 5
@@ 5.1.1 PH-based approaches are robust to zero-inflation
We simulate spatial data as described in Section 4.5 and first vary the zero-inflation parameter $z$ from $0.1$ to $0.99$, keeping the dispersion fixed at $0.3$. For each shape and parameter choice, we then simulate $50$ genes that possess the spatial signal and $50$ genes that are randomly sampled without any spatial effect. Then, across each degree of zero-inflation, we summarise the performance of each statistical test and scoring method in the following manner: Fig. 3 reports the area under the precision-recall curve (AUPRC) as a function of the degree of zero-inflation for all methods. This comparison is chosen because it allows us to compare all methods, also the ranking-based sepal method, as a single summary statistic per degree of zero-inflation. An AUPRC of $1$ indicates perfect performance, whereas random guessing is expected to result in an AUPRC of $0.5$, and 00 is the worst possible value.
@@ 5.1.2 Persistence-based testing performs well at detecting SVGs
We observe that all persistence-based methods achieve high AUPRC values very close to $1$ until around $70\%$ of zero-inflation across all patterns, as summarised in Fig. 3. This is followed by a decrease in AUPRC, which remains above $0.6$, even for $99\%$ excess zeros. Overall, these trends show the robustness and expressivity of persistent homology as a summary of spatial patterns, even in the presence of high zero-inflation.
@@ 5.1.3 Persistence-based methods perform similarly to one another
We further observe that both Betti curves and total persistence follow very similar trends in Fig. 3, with Betti curves performing better on the gradient pattern and total persistence showing higher scores on the streaks pattern. This finding highlights that both approaches are alike in terms of the information they are summarising. Persistent landscapes meanwhile achieve consistently higher AUPRC values above $0.7$ for higher degrees of zero-inflation across datasets and thus capture trends in spatial variability in a more robust manner for high degrees of dropout. For this experiment, we thus observe that persistence landscapes summarise different aspects of persistent homology than alternative summaries. This slight difference in performance reported in Fig. 3 arises because persistence landscape functions [9] differ in the way they are constructed from a given persistence diagram compared to Betti curves, which are more alike to total persistence [37].
5.1.4 Persistence-based methods outperform Moran’s I on zero-inflated data
Our results in Fig. 3 indicate that persistent homology is more expressive at quantifying spatial signals in simulated expression values under the presence of zero-inflation than Moran’s I. That is, we observe that by tracking the evolution of connected components via superlevel-set filtrations on spatial graphs, we get summary statistics with higher discriminative performance than Moran’s I, which summarises spatial autocorrelation based on summarising local neighbourhoods and is the most commonly used baseline measure for spatial variability. Indeed, via this comparison in terms of AUPRC, we assess that persistence-based summaries offer a valid alternative to established tools from spatial statistics.
@@ 5.1.5 Persistence detects spatial variability across all simulations
We note that our proposed persistence-based tests do well at detecting spatial variability, in particular for low levels of zero-inflation, as would be expected. However, in comparison we see that some alternative methods do not follow the same clear trend across datasets. SpatialDE starts decreasing in terms of AUPRC at comparably lower proportions of zero inflation. This difference is especially pronounced for the streaks pattern, where SpatialDE performs notably worse than all alternative methods after $60\%$ zero-inflation. Potentially this is due to this simulation only showing two relatively thin lines (as plotted in Fig. 2) that are harder for SpatialDE to detect as spatially variable under the presence of excess zeroes. Meanwhile, sepal performs very well for higher values of $z$ between $0.6$ and $0.8$ across datasets, showing higher AUPRCs than persistence landscapes. However, sepal fails at detecting the gradient pattern at low degrees of zero-inflation. For example, sepal receiving an AUPRC of only $0.6$ at $z=0.1$ highlights that diffusion times are not suitable for detecting this smooth trend in spatial variability. We thus demonstrate disadvantages of both sepal and SpatialDE compared to our persistence-based approach. We conclude that PH-based methods achieve overwhelmingly higher or similar AUPRCs to the baseline method Moran’s I, highlighting their utility as robust SVG detection methods.
@@ 5.1.6 PH-based methods are more specific than Moran’s I
Extending our observations beyond AUPRC, we next decide to view spatially variable gene detection in the context of hypothesis testing. Fig. 4 compares the performance of each statistical test in terms of specificity and sensitivity at a fixed $0.05$ significance threshold for corrected p-values. In agreement with results from Byers et al. [11], we observe that at this fixed threshold, total persistence is either similarly or slightly less powerful, i.e. less sensitive than Moran’s I. This is followed by Betti curves, which detect a lower proportion of true positives, in particular on the streaks pattern, as well as persistence landscapes which give even more conservative estimate. However, we also observe that all persistence-based approaches show very high specificity across examples, whereas Moran’s I demonstrates the highest false positive rate of around $5\%$ to $25\%$ even after p-value correction, which is undesirable in practice when the set of spatially variable genes should be correctly identified.
Our findings thus agree with Byers et al. [11], who find that persistence is less sensitive than Moran’s I. However, Byers et al. [11] further claim that persistence cannot replace Moran’s I, a finding we do not fully support. Arguably, specificity as well as sensitivity are important for interpreting spatial dependency tests in practice. Indeed, we observe that at a default $0.05$ significance cutoff, corrected p-values from PH-based approaches achieve higher specificity than Moran’s I, showing that PH-based results can be more useful for tasks that require a low false positive rate and benefit from a more conservative assessment. Given the high AUPRC achieved by PH-based methods, we further assess that choosing a higher significance threshold could be a suitable adjustment to achieve higher sensitivity and identify a larger set of spatially variable features while retaining superior specificity.
@@ 5.1.7 Agreement between methods
Figure 5: Spearman correlation between the rankings given by different spatial variability detection methods across the four spatial patterns from Fig. 2. Generally, there is low to medium correlation between methods with the highest agreement being between total persistence and Betti curves. Persistence thus captures distinct characteristics in spatial variability not identified by other approaches.
Further, we compute the Spearman correlation between each method’s ranking and report the results in Fig. 5. For persistence-based approaches, we rank genes by their p-values, else we use the values of the scores as defined in Section 4.3 for sepal and Moran’s I. From the results in Fig. 5, we observe that total persistence and Betti curves perform most similarly across experiments with correlations between 0.64 and 0.82, which is likely driven by their common roots in accumulating persistence features [37]. Further, we find that there is varying but overall moderate levels of agreement between persistence and Moran’s I ($\rho\leq 0.53$), likely explained by the fact that both are summaries of the same spatial graph and understand spatial variability via modelling clustering behaviour, as summarised by connected components in PH, or autocorrelation, as summarised by differences in neighbourhood scores. Interestingly, while sepal follows a fundamentally different modelling approach, correlation between sepal and the other approaches exists at comparable levels to other entries in the correlation matrix and the sepal score is actually most correlated with Moran’s I on the clusters dataset ($\rho=0.58$). Overall, we thus find that persistence-based permutation testing captures complementary information not encoded by either Moran’s I, sepal, or SpatialDE.
@@ 5.2 PH-based approaches identify the top SVGs
Next, we keep the zero proportion at $z=0.1$ and the dispersion parameter at $r=0.3$, following estimations by Zhu et al. [47] from real data, but vary the effect size by taking $\hat{e_{i}}\leavevmode\nobreak\ =\leavevmode\nobreak\ max(e_{i}/c,1)$ for some $c$ in $\{6,5,4,3,2,1\}$ for each domain $i=1,2,...,N$ effectively varying the strength of the spatial pattern compared to the background domain, for which $e_{0}=1$. For each parameter choice and spatial shape, we then simulate 50 features with spatial signals and 50 completely random patterns. We then investigate the identification of top spatially variable genes for different strengths of spatial effects.
@@ 5.2.1 Correlation with the sum of gene expression counts
As reported in Fig. 6, we explore the correlation between each spatial variability score and the total count per gene, i.e. the sum of feature values. We want an ideal SVG detection method to not just correspond to these total counts, but score genes based solely on the existence of spatial dependencies in gene expression. Across datasets, we see that Moran’s I and sepal show the highest correspondence to the sum of feature values. In comparison, persistence-based rankings achieve consistently lower Spearman correlation of $0.6$ to the number of total counts. Persistence landscapes reach slightly higher correlation than Betti curves and total persistence. Overall, this rank correlation is similar to the one achieved by SpatialDE. Therefore, the comparison highlights that persistence offers a satisfactory alternative to established SVG tests.
@@ 5.2.2 PH-based approaches correctly identify the top SVGs
Across methods, we then compare the ground truth given by the existence of spatial signal with each spatial variability score. Fig. 6 shows the proportion of truly positive features detected by each method. Ideally, a suitable spatial variability detection method would show a high proportion of true positives amongst the top 300 genes (given that 300 out of 600 features show spatial patterns). Notably, Betti curves perform consistently well across examples correctly identifying more than $90\%$ of top SVGs. This is followed by total persistence and persistence landscapes, which perform better on the gradient but worse on the cluster pattern. Interestingly, from these examples we see that while the proportion of genes correctly identified is high across scenarios, there is no strong distinction between either persistence- or non-persistence-based methods based on these examples. However, sepal and Moran’s I seem to perform slightly worse at identifying the most spatially variable features than persistence-based approaches, which together with their correlation to total counts indicates that persistence-based testing is preferable in this scenario.
@@
Figure 9: Examples of the most spatially variable genes as identified by Betti curves for the mouse olfactory bulb (top), the breast cancer sample (middle), and the heart data (bottom). Each plot shows the spatial distribution of one gene’s expression values preprocessed as detailed in Section 4.4. The name of each gene is plotted above the respective tissue sample.